OpenStudy (anonymous):

An ecology center wants to set up an experimental garden using 300 m of fencing to enclose a rectangular area of 5000 m2. Find the dimensions of the garden. fence length 150-x and width equals x

OpenStudy (cwrw238):

area = x(150 - x) = 5000 solve for x

OpenStudy (anonymous):

150x-x2=5000 150x-x2-5000=0 is this right so far?

OpenStudy (cwrw238):

yes but write it as x^2 - 150x + 5000 = 0

OpenStudy (cwrw238):

(x - 50)(x - 100) = 0 x = 50 or 100 so dimensions of the grade is 50 by 100 m.